# Dreaming of Many Worlds: Learning Contextual World Models Aids Zero-Shot Generalization

### v1

@@ Acknowledgments
Sai Prasana and Karim Farid acknowledge funding by the Konrad Zuse School of Excellence in Learning and Intelligent Systems (ELIZA) grant. Raghu Rajan and André Biedenkapp acknowledge funding from The Carl Zeiss Foundation through the research network “Responsive and Scalable Learning for Robots Assisting Humans” (ReScaLe) of the University of Freiburg.
@@ Appendix B Train and Evaluation Context Ranges
Context |
Gravity |
Length |
Default |
9.8 |
.5 |
Training Range |
[4.9, 14.7] |
[.35, .75] |
Single Evaluation Values |
.98, 17.15, 2.45, 3.92, 4.9, 7.35, 9.8, 12.25, 14.7, 15.68, 16.66, 17.64, 18.62, 19.6 |
.1, .15, .2, .25, .3, .4, .5, .6, .7, .8, .85, .9, .95, 1.0 |
Dual Evaluation Values |
.98, 2.45, 3.92, 15.68, 17.64, 19.6 |
.1, .2, .3, .5, .7, .8, .9, 1.0 |
Context |
Gravity |
Actuator Strength |
Default |
9.8 |
.5 |
Training Range |
[4.9, 14.7] |
[.5, 1.5] |
Single Evaluation Values |
.98, 17.15, 2.45, 3.92, 4.9, 7.35, 9.8, 12.25, 14.7, 15.68, 16.66, 17.64, 18.62, 19.6 |
.1, .2, .3, .4, .5, .75, 1.0, 1.25, 1.5, 1.6, 1.7, 1.8, 1.9, 2.0 |
Dual Evaluation Values |
.98, 2.45, 3.92, 15.68, 17.64, 19.6 |
.1, .3, .5, 1.0, 1.5, 1.6, 1.8, 2.0 |

### v2

@@ Acknowledgments
S.P. and K.F. acknowledge funding by the Konrad Zuse School of Excellence in Learning and Intelligent Systems (ELIZA) grant. R.R. and A.B. acknowledge funding from the Carl Zeiss Foundation through the research network “Responsive and Scalable Learning for Robots Assisting Humans” (ReScaLe) of the University of Freiburg.
@@ Appendix B Train and Evaluation Context Ranges
Context | Gravity | Length |
Default | 9.8 | .5 |
Training Range | [4.9, 14.7] | [.35, .75] |
Single Evaluation Values | .98, 17.15, 2.45, 3.92, 4.9, 7.35, 9.8, 12.25, 14.7, 15.68, 16.66, 17.64, 18.62, 19.6 | .1, .15, .2, .25, .3, .4, .5, .6, .7, .8, .85, .9, .95, 1.0 |
Dual Evaluation Values | .98, 2.45, 3.92, 15.68, 17.64, 19.6 | .1, .2, .3, .5, .7, .8, .9, 1.0 |
Context | Gravity | Actuator Strength |
Default | 9.8 | .5 |
Training Range | [4.9, 14.7] | [.5, 1.5] |
Single Evaluation Values | .98, 17.15, 2.45, 3.92, 4.9, 7.35, 9.8, 12.25, 14.7, 15.68, 16.66, 17.64, 18.62, 19.6 | .1, .2, .3, .4, .5, .75, 1.0, 1.25, 1.5, 1.6, 1.7, 1.8, 1.9, 2.0 |
Dual Evaluation Values | .98, 2.45, 3.92, 15.68, 17.64, 19.6 | .1, .3, .5, 1.0, 1.5, 1.6, 1.8, 2.0 |
@@ Appendix F Results for DMC Walker - 100k Steps
Appendix G Discussion on More Comprehensive Benchmarks
Creating contextual benchmarks for environments such as Atari, DMLab, ProcGen, and Minecraft presents an important opportunity for further research into ZSG. Unlike our current tasks focusing on motor control in environments like Walker and CartPole, some of these benchmarks emphasize different aspects such as navigation and exploration while others such as procedurally generated or open-ended worlds offer dynamic objectives and high variability, requiring strategic planning and adaptability to diverse challenges. These complexities necessitate pronounced policy adjustments with changing contexts. However, creating variants in such benchmarks is challenging due to their inherent intricacies and lack of easily accessible interfaces. By addressing these complexities, future work could unlock critical insights into the adaptability of reinforcement learning agents in diverse and changing conditions.